OpenStudy (anonymous):

is there a way to simplify 1/cos^2theta*tan^2theta? or how would you simplify (sectheta*cottheta)^2?

OpenStudy (anonymous):

\[1/\cos^2\theta \times \tan^2 \theta \]

OpenStudy (anonymous):

like that?